# Automated Testing Method for Text-to-Image Software

### v2

@@ 3.1. Adaptability Density-guided Mutation Operators
We note that not all words in the input text have an impact on the quality of the results generated by T2I software. Therefore, traditional text mutation operators are not effective in thoroughly testing the defects in T2I software. As we mentioned in Sec 2.3, mutation operators should be highly relevant to focal ERs across two modalities. The problem we need to address is how to guide and design these mutation operators effectively and then generate the testing samples.
In the T2I software testing phase, items can be characterized by a set of exercises $M$. Testers use a certain measurement $R$ to test each exercise $\mu\in M$. The performance of software $\pi$ under an exercise $\mu$ can be denoted as $R(\pi,\mu)$. Because the measurements often need to be made multiple times, it is usual to apply the expected value of the performance of $\pi$ as $E[R(\pi,\mu)]$. As the T2I software structure is complicated, $M$ may contain more than one problem. The metrics $\Phi$ most commonly used to describe aggregated results are:
@@
where $p(\mu)$ is the probability of $\mu$. T2I software $\pi$ is large and complicated enough that we clearly cannot test $\pi$ on the whole set $M$. Random sampling with $p$ is a reasonable choice. If $M$ and p define the benchmark, is probability-proportional sampling on $p$ the best way to test software? The answer is no, in general. Because the set $M$ can include very easy and very challenging problems simultaneously (e.g. text summarization, text understanding, question answering, sentiment analysis, relation extraction, image generation, image super-resolution, and image composition). $R$ can also be nondeterministic and/or subject to measurement error. Therefore, the probability-proportional sampling can be inefficient and costly. The better idea to approximate Eq.1 is to sample in a more purposeful way.
Our testing objective is to test the generation robustness of the T2I software. Robustness denotes the ability of software to consistently operate across diverse situations and the performance of the software to abnormal inputs and conditions. Essentially, we seek to understand the software’s adaptability density with respect to varied test samples. In the example of Figure 3, the fourth image illustrates the highest adaptability density compared with others because it successfully synthesizes all the entities (”man”, ”bird”, ”stone”) and relationships (”watching”, ”standing on”).
Therefore, to better optimize the sampling method in the testing phase, we need to define the adaptability density constraint $d:M\rightarrow\mathbb{R}^{+}$. In the most ideal state, for each $\pi$ being tested, $\Phi\left(\pi,M_{1},p\right)>\Phi\left(\pi,M_{2},p\right)\ if\ d\left(\mu_{1}\right)<d(\mu_{2})$. That is, the higher the adaptability density is, the better the evaluation performance is. As a result, we formulate the adaptability density-driven sampling method as follows:
(2) $\Phi\left(\pi,M,h\right)=\ \sum_{\mu\in M}\Phi\left(\pi,M,h\right)=\sum_{\mu\in M}{h(\mu)\bullet p(\mu|h(\mu))}\bullet\mathbb{E}[R(\pi,\mu)]$
where $h(\mu)$ is the probability of $\mu$. For better instantiation and a more flexible condition, we discretize $h$. The adaptability density constraint is finally improved to that for each $\pi$ being tested and two adaptability density levels $a$ and $b$ ($a\leq b$), we can get that $\Phi\left(\pi,M_{a},p\right)>\Phi\left(\pi,M_{b},p\right)$ (where ${M}_{a}={\mu|\mu\in M,d\left(\mu\right)=a}$).
@@
Typically, $\pi$ remains a black box to end-users, obscuring the exact techniques it employs. However, it is unequivocally necessary for this task to align text and images, implying a requirement for consistency in entities and their interrelationships, albeit represented differently. Given these considerations, we define the adaptability density constraint corresponding to the matching degree of semantic information in two modalities.
@@
However, it is difficult to directly calculate the adaptability density of T2I software based on the semantic information. The contrastive language-image pre-training (CLIP) (Radford et al., 2021) model is not suitable for this purpose, as it is highly vulnerable to attacks and lacks sufficient interpretability to provide further explanations to testers for anomalies. Furthermore, using CLIP as a step in the testing process is unfair because some T2I software (e.g. DALLE) utilizes this model as a pre-trained model. Therefore, to better represent semantic information in cross-modal data, and given that the input text for T2I software is mostly scene descriptions, which is highly analogous to scene understanding in the field of computer vision, we adopt the ER model to represent semantic information.
@@
As shown in Figure 1, we deliver four output images corresponding to two input texts. The salient entities in each image are consistent with the input texts, marked in red (e.g. ”man”, ”motorcycle”, ”cat”). If the input text describes the relationship between focal entities, marked in green (e.g. ”standing next to”, ”and”, ”sleep on”), the generated image should describe as well. Each entity and relationship has its class. Based on the entity and relation extraction technique in nature language processing, each text can be constructed to an entity-relation pool (ER pool), containing several entity-relation triples (e.g. (”dog”, ”with”, ”cat”), (”dog”, ”on”, ”bed”), (”cat”, ”on”, ”bed”)). All elements in the ER pool will be converted to the class they belong to (e.g. ”and” convert to ”with”, ”sleep on” convert to ”on”, ”man” convert to ”person”) referring to (Chiou et al., 2021; Yu et al., 2023). The focus of the paper does not include the analysis of the singular and plural nouns.
The absence of test oracles in the T2I software makes it necessary to apply the metamorphic testing method. Metamorphic testing is introduced as a solution to the test oracle problem in test case generation (Chen et al., 2018). It enables the generation of transformed tests from successful ones and mitigates potential issues arising from the lack of test oracles by defining the concept of Metamorphic Relations (MR). MR represents an essential characteristic of the target function or algorithm concerning various inputs and their corresponding expected outputs. Hence, the breach of MR serves as an indicator of potential software defects (Segura et al., 2016). Establishing a suitable MR forms the foundation for conducting metamorphic testing.
Formally, we denote the T2I software as ${f_{T2I}}$, the input text in the seed set $\mathbb{S}=\{s_{i}\}$. Due to the unique nature of generative tasks, we do not use the existing label. We define the testing method in ACTesting as $f_{er}$, which is used to measure the degree of cross-modal semantic information matching. The results can be formalized as follows:
(3) $r_{s_{i}}={f_{er}}(f_{T2I}(s_{i}))$
Then we define the transformation operators $f_{m}$ to generate a set of new text $\mathbb{A}=\{a_{j}\}$. The mutation results (generated image) can be formalized as follows:
(4) $r_{a_{j}}={f_{er}}(f_{T2I}(f_{m}(s_{i})))$
The testing domain is defined for each generation result as follows:
(5) $\mathbb{M}_{s_{i}}=\{m_{j}\|d(n_{s_{i}},m_{j})<\epsilon\}$
@@
where $\epsilon$ is a parameter for measuring strategy. Finally, the MR to test the T2I software with the newly generated sample can be formalized as follows:
(6) $\forall s_{i}\in\mathbb{S}\wedge\forall a_{j}\in\mathbb{A},r_{a_{j}}\in\mathbb{M}_{s_{i}}$
Different from the traditional text mutation operators, ACTesting is aimed at performing robustness testing for T2I software. As mentioned in Sec 2.3, we observe that when the same sentence undergoes a change in a single entity, there are differences in the generated quality exhibited by the T2I software. Additionally, as the number of entities and relationships in the sentence increases, the software’s processing capabilities also show some difference. Therefore, based on the adaptability density constraint we introduced, three kinds of transformation operators are designed. We do not directly mutate the input text; instead, we first construct each text into an ER pool. Figure 5 presents an example of implementation for mutation operators of ACTesting.
@@
Entity Changing (EC): The EC operator substitutes one entity in the ER pool, thus constructing a pair of highly similar input text samples. The output image should ideally remain consistent, aside from the substituted entity after generation. EC operator is primarily employed to ascertain whether T2I software demonstrates variability in generating content for diverse entities. In essence, the EC operator assesses the adaptability of T2I software, examining whether it excessively relies on pre-existing ER in the training dataset, consequently resulting in constrained generation capabilities. EC operator focuses on the issue of generation stability within the context of robustness testing.
@@
Entity-Relationship Removal (ER_R): The ER_R operator randomly removes an ER triple. Removing one ER triple does not lead to unreadable sentences or semantic errors, which sets it apart from traditional text mutation strategies. The generation of images should ensure that the retained entities and relationships remain unaffected and maintain consistent image quality. ER_R operator aims at constructing a pair of text samples with differing levels of complexity, used to investigate whether the T2I models exhibit significant disparities in generating these different samples, which is also the reflection of the aspect of generating robustness.
@@
Entity-Relationship Augmentation (ER_A): The ER_A operator randomly adds an entity and its corresponding relationship from candidate entities and relationships to the original ER pool. We construct the candidate ERs from the dataset. After obtaining the new ER pool, we use it as a basis to rebuild the input text. By increasing the variation of the original text input, ACTesting can explore the search space more extensively to uncover potential issues. ER_A operator aims to increase the complexity of the original samples, expanding the input space to test the robustness of T2I software.
Figure 5. The mutation operators of ACTesting
In summary, the design of the three mutation operators is aimed at detecting defects in T2I software, based on mutation relationships that maintain consistency with cross-modal semantic information.
3.2. ACTesting Workflow
To ensure the adaptability of density-guided mutation operators can test the robustness of T2I software, ACTesting also encompasses the construction of a complete testing process. As shown in Figure.4, ACTesting includes several main processing components: ER pool construction, adaptability density-guided operators, text synthesizing, image entity and relationship detection, and testing and evaluation results calculation. We select representative seed texts from the dataset that contain diverse ERs at the beginning. Then, ACTesting constructs an ER pool for each seed text based on the principle of semantic information preservation.
ER Pool Construction. The input text in the T2I software is different from other natural language processing tasks like text Classification or text abstract. The text often contains the scene or the objects of the generated image, which renders traditional text mutation operators no longer applicable to inputs for this task. Such as swapping the order of the words or introducing spelling errors in the text. Considering the different ways in which various modalities express information, we extract each input text as an ER pool, utilizing consistent cross-modal semantic information features. Subsequently, all mutation operations are performed based on the ER pool, ensuring that the semantic features of the input text remain intact. We adopt the relationship extraction, entity recognition and structured representation technique to construct the ER pool referring (Qi et al., 2020). The ER candidates are also construed by this method based on the open-source dataset.
Mutation Operator Implementation As introduced in Sec 3.1, three kinds of mutation operators are aimed at testing the generation robustness of T2I software through cross-modal alignment. ACTesting adopts the EC and ER_A operators to all the inputs and adopts the ER_R operator to the input having more than one ER. The transformation operators of ACTesting are shown in Fig.5.
Text Synthesizing. After completing the mutation operations based on the adaptability density-guided operators, ACTesting constructs the synthesized text based on the mutated ER pools. To ensure the fluency of synthesized sentences and the integrity of semantic information, we define templates for generating sentence structures and utilize BERT (Devlin et al., 2018) to enhance the details of the sentences. At this point, we have obtained paired text inputs.
Image object and relationship detection. We send the paired text inputs to the tested T2I software, respectively, and generate the output images. To address the challenge of cross-modal semantic matching, we employ the object detection model to identify entities in images and use the scene understanding model to detect relationships between entities in images. In the end, we forward the detected results to the next component.
Testing and Analysis. We consider that the degree of matching between images and text is fundamentally based on the matching degree of entities and relationships. Based on the selection of previous mutation operators, different MRs are used to compute test results. The formalized MR is introduced in the Sec 3.1. Specifically, we design MRs for each mutation operator as follows:
(7) $K_{e}=Set_{e}(r_{s_{i}}))\cap Set_{e}(r_{a_{i}}))$
(8) $K_{r}=Set_{r}(r_{s_{i}}))\cap Set_{r}(r_{a_{i}}))$
where $Set_{e}(r_{s_{i}}))$ and $Set_{e}(r_{a_{i}}))$ denote the set of entity class in $r_{s_{i}}$ and $r_{a_{i}}$, $Set_{r}(r_{s_{i}}))$ and $Set_{r}(r_{a_{i}}))$ denote the set of relationship class in $r_{s_{i}}$ and $r_{a_{i}}$.
MR1 is designed for the EC operator:
(9) $K_{e}==Set_{e}$ $(r_{s_{i}}))-Set_{e}(e_{1})==Set_{e}(r_{a_{i}}))-Set_{e}(e_{2})$
(10) $K_{r}==Set_{r}(r_{s_{i}}))==Set_{r}(r_{a_{i}}))$
MR2 is designed for the ER_R operator:
(11) $K_{e}==Set_{e}(r_{a_{i}}))$
(12) $K_{r}==Set_{r}(r_{a_{i}}))$
MR3 is designed for the ER_A operator:
(13) $K_{e}==Set_{e}(r_{s_{i}}))$
(14) $K_{r}==Set_{r}(r_{s_{i}}))$
@@ 3.2. ACTesting Workflow
If any of the $K_{e}$ or $K_{r}$ in each MR are violated under a certain operator, the generated results will be separately reported as either entity errors ($p_{e}$) or relationship errors ($p_{r}$).
In summary, we detail the specific implementation of metamorphic relationships tailored for testing T2I software. We will delve into the specific evaluation metrics related to that in the following section. ACTesting encompasses a comprehensive process of test case generation and test result computation, where each module plays a crucial role in the overall effectiveness of the testing.

### v3

@@ 3.2. Adaptability Density-guided Constraint
As we mentioned in Sec 2.3, changes in the focal ERs in the input text can easily affect the generated results of T2I software. This means that not all words in the input text impact the quality of the results generated by T2I software. Therefore, traditional random text mutation operators are not effective in thoroughly testing the errors in T2I software. To address this problem, we explore the adaptability density-guided constraint to guide the mutation operators to generate the new input text effectively.
We first formalize the testing process for the T2I software to better understand our proposal. In the T2I software testing phase, one test can be characterized by a set of scenarios $M$ which represent specific input configurations, such as individual text inputs. Each test scenario ${\mu\in M}$ corresponds to a distinct configuration and is evaluated using a performance measure $R$. The performance of software $\pi$ under a given test scenario $\mu$ can be denoted as $R(\pi,\mu)$. To account for variations caused by non-deterministic behaviors or random factors, the expected performance $E[R(\pi,\mu)]$ is utilized, representing the average outcome over multiple test runs for scenario $\mu$. Given the complexity of T2I software structures, different scenarios often yield varying performance results. The overall performance of the software across all scenarios in $M$ is quantified using the aggregated metric $\Phi\left(\pi,M,p\right)$. This metric incorporates $p(\mu)$, which denotes the probability or normalized weight of executing scenario $\mu$, to reflect its relative importance. The aggregated performance metric is formally defined as:
@@
T2I software $\pi$ is large and complicated enough which cannot be tested across the entire set $M$. Random sampling guided by the distribution $p$ is often employed. However, when $M$ and $p$ are used to define the testing benchmark, probability-proportional sampling based on $p$ may not always provide the most effective testing strategy. Because the set $M$ can include different problems simultaneously (e.g., text understanding, image generation, and image composition), each of which may require different testing priorities. $R$ can also be nondeterministic and/or subject to measurement error. Therefore, the probability-proportional sampling can be inefficient and costly.
Given the inefficiencies and potential costs of probability-proportional sampling, a more purposeful approach to sampling is necessary to effectively approximate Eq. 1. ACTesting focuses on testing the generation quality of T2I software, which is defined as the software’s ability to perform correctly across diverse scenarios. To achieve this, we aim to assess the software’s adaptability density—its capacity to handle varied test inputs while maintaining functionality. For instance, as shown in Fig. 3, Image 2 illustrates the higher adaptability density compared with Image 1 because it successfully synthesizes all the entities (“man”, “bird”, “stone”) and relationships (“watching”, “standing on”).
To optimize the sampling method in the testing phase, we introduce the adaptability density constraint $d:M\to\mathbb{R}^{+}$. This constraint quantifies the adaptability density of each test scenario $\mu\in M$. Ideally, for any T2I software $\pi$ being tested, the following relationship holds:
$\Phi\left(\pi,\mu_{1},p\right)>\Phi\left(\pi,\mu_{2},p\right)\quad\text{if}\quad d\left(\mu_{1}\right)<d\left(\mu_{2}\right).$
In other words, higher adaptability density $d(\mu)$ corresponds to improved evaluation performance. Therefore, we propose the adaptability density-driven sampling method, formulated as:
$\Phi\left(\pi,M,h\right)=\sum_{\mu\in M}\Phi\left(\pi,\mu,h\right)=\sum_{\mu\in M}h(\mu)\cdot p(\mu\mid h(\mu))\cdot\mathbb{E}[R(\pi,\mu)],$
where $h(\mu)$ denotes the sampling probability of scenario $\mu$, weighted by its adaptability density. Under the T2I task, $h$ is discretized to better account for varying adaptability density levels. To refine the adaptability density constraint, we assert that for any T2I software $\pi$ being tested and for two adaptability density levels $a$ and $b$ ($a\leq b$), the following relationship is satisfied:
$D(\Phi\left(\pi,M_{a},h\right),\Phi\left(\pi,M_{b},h\right))<\epsilon,$
where $D(\cdot,\cdot)$ is a distance metric applied to the evaluation results $\Phi$ aggregated over the respective scenario subsets ($M_{a}$ $M_{a}=\{\mu\mid\mu\in M,d(\mu)=a\}$) and $M_{b}$ ($M_{b}=\{\mu\mid\mu\in M,d(\mu)=b\}$). $\epsilon$ is a parameter of constraint. This ensures that scenarios with lower adaptability density contribute more significantly to the evaluation, promoting a robust and balanced assessment of the software’s performance.
@@
Typically, $\pi$ remains a black box to end-users, obscuring the exact techniques it employs. However, it is unequivocally necessary for this testing task to align text and images, implying a requirement for consistency in entities and their interrelationships, albeit represented differently. Given these considerations, we specify the adaptability density constraint corresponding to the matching degree of semantic information in two modalities.
@@
However, it is difficult to calculate the adaptability density of T2I software directly based on the semantic information. The contrastive language-image pre-training (CLIP) (Radford et al., 2021) model is unsuitable, as it is highly vulnerable to attacks and lacks sufficient interpretability to provide further explanations to testers for anomalies. Therefore, to better represent semantic information in cross-modal data, and given that the input text for T2I software is mostly scene descriptions, which is highly analogous to scene understanding in the field of computer vision, we adopt the ER model to represent semantic information.
@@
Specifically, as shown in Fig. 1, we deliver two output images corresponding to two input texts. The salient entities in each image are consistent with the input texts, marked in red (e.g., “man”, “motorcycle”, “cat”). If the input text describes the relationship between focal entities, marked in green (e.g., “standing next to”, “and”, “sleep on”), the generated image should be described as well. Each entity and relationship has its class. Based on the entity and relation extraction technique in nature language processing, each text can be constructed to an entity-relation pool (ER pool), containing several entity-relation triples (e.g., (“dog”, “with”, “cat”), (“dog”, “on”, “bed”), (“cat”, “on”, “bed”)). All elements in the ER pool are converted to the class they belong to (e.g., “and” convert to “with”, “sleep on” convert to “on”, “man” convert to “person”) referring to (Chiou et al., 2021; Yu et al., 2023). Note that the analysis of singular and plural nouns is not the focus of our proposal.
3.3. Metamorphic Relation
One of the central challenges in testing T2I software is the oracle problem—the difficulty of objectively determining whether the output matches the target. This challenge arises from the inherent subjectivity in assessing complex outputs like images, where the correspondence between input prompts and generated visuals cannot be easily quantified. To address these limitations, we adopt MT and define the concept of MR to mitigate potential issues arising from the lack of test oracles. As we analyze in Sec 3.2, the focal entities and relationships contained in the input text and output image are the primary semantic representations. Therefore, ACTesting uses the ER model to represent the cross-model semantic and takes semantic consistency as the basic MR.
Formally, considering the T2I software as $\pi$, the input text in the seed set $\mathbb{S}=\{s_{i}\}$ ($s_{i}$ represents the $i$-th input text). The adaptability density-guided constraint in ACTesting is defined as $d$. $f_{er}$ is used to represent the cross-modal semantic information. The result obtained after the output image generated by the T2I software based on the input $s_{i}$ is characterized by the ER model can be represented as:
(2) $r_{s_{i}}={f_{er}}(\pi,s_{i})$
Then, we construct new text ($\mathbb{A}=\{a_{j}\}$ ($a_{j}$ represents the $j$-th input text)) based on the adaptability density constraint $d$. Similarly, we use Eq 2 to obtain the result characterized by the ER model:
(3) $r_{a_{j}}={f_{er}}(\pi,f_{m}(s_{i}))$
@@ 3.3. Metamorphic Relation
After obtaining the two outputs, we apply the MR to determine whether the generation requirements of the T2I software are met. Finally, the MR to test the T2I software with the newly generated sample can be formalized as follows:
(4) $\forall s_{i}\in\mathbb{S},\ \forall a_{j}\in\mathbb{A},D(r_{s_{i}},r_{a_{j}})<\epsilon$
where $\epsilon$ represents the parameter of the constraint.
Figure 5. The examples of applying mutation operators of ACTesting.
3.4. Mutation Operators
Based on the context of Sec 3.2 and Sec 3.3, it is crucial to generate new input samples guided by the adaptability density constraint. To this end, we introduce three types of mutation operators designed around the proposed adaptability density constraint. In the construction of text inputs, entities are central elements and are more likely to trigger generation errors in T2I software. Relationships, on the other hand, often serve as subordinate or linking components for entities. Therefore, when designing mutation operators guided by the adaptability density constraint, we define separate principles for entity changing and the unified augmentation or removal of ERs. This ensures semantic validity and preserves the integrity of core content. Fig. 5 illustrates an example of the implementation of these mutation operators in ACTesting. To specify the MR based on cross-modal semantic consistency for T2I software, we refer to  (Yu et al., 2022b) and define the MR and corresponding mutation operators as follows:
(5) $K_{e}=Set_{e}(r_{s_{i}})\cap Set_{e}(r_{a_{i}})$
(6) $K_{r}=Set_{r}(r_{s_{i}})\cap Set_{r}(r_{a_{i}})$
where $Set_{e}(r_{s_{i}})$ and $Set_{e}(r_{a_{i}})$ denote the set of entity class in $r_{s_{i}}$ and $r_{a_{i}}$, $Set_{r}(r_{s_{i}})$ and $Set_{r}(r_{a_{i}})$ denote the set of relationship class in $r_{s_{i}}$ and $r_{a_{i}}$.
@@ 3.4. Mutation Operators
Entity Changing (EC): The EC operator substitutes one entity in the ER pool, thus constructing a pair of highly similar input text samples. The output image should ideally remain consistent, aside from the substituted entity after generation. EC operator is primarily employed to ascertain whether T2I software demonstrates variability in generating content for diverse entities. In essence, the EC operator assesses the adaptability of T2I software, examining whether it excessively relies on pre-existing ER in the training dataset, consequently resulting in constrained generation capabilities. EC operator focuses on the issue of generation stability within the context of testing. MR detailed for the EC operator:
(7) $K_{e}==Set_{e}$ $(r_{s_{i}})-Set_{e}(e_{1}),K_{e}==Set_{e}(r_{a_{i}})-Set_{e}(e_{2})$
(8) $K_{r}==Set_{r}(r_{s_{i}}),K_{r}==Set_{r}(r_{a_{i}})$
@@
Entity-relationship Removal (ER_R): The ER_R operator randomly removes an ER triple. Removing one ER triple does not lead to unreadable sentences or semantic errors, which sets it apart from traditional text mutation strategies. The generation of images should ensure that the retained entities and relationships remain unaffected and maintain consistent image quality. ER_R operator aims at constructing a pair of text samples with differing levels of complexity, used to investigate whether the T2I models exhibit significant disparities in generating these different samples, which is also the reflection of the aspect of generating quality. MR detailed for the ER_R operator:
(9) $K_{e}==Set_{e}(r_{a_{i}})$
(10) $K_{r}==Set_{r}(r_{a_{i}})$
@@
Entity-relationship Augmentation (ER_A): The ER_A operator randomly adds an entity and its corresponding relationship from candidate entities and relationships to the original ER pool. We construct the candidate ERs from the dataset. After obtaining the new ER pool, we use it as a basis to rebuild the input text. By increasing the variation of the original text input, ACTesting can explore the search space more extensively to uncover potential issues. ER_A operator aims to increase the complexity of the original samples, expanding the input space to test the generation quality of T2I software. MR detailed for the ER_A operator:
(11) $K_{e}==Set_{e}(r_{s_{i}})$
(12) $K_{r}==Set_{r}(r_{s_{i}})$
@@
If any of the $K_{e}$ or $K_{r}$ in MR are violated under a certain operator, the generated results will be separately reported as either entity errors ($p_{e}$) or relationship errors ($p_{r}$). In summary, the design of the three mutation operators aims to detect errors in T2I software based on MR that maintain consistency in cross-modal semantic information.